# Logistic Multidimensional Data Analysis for Ordinal Response Variables using a Cumulative Link function

### v1

@@ 4 Application
4.1 Cumulative Logistic Reduced Rank Regression
In this analysis, we take the four behavioural items as response variables. Predictors are environmental concern (treated as a numeric variable) and environmental efficacy which we obtained by taking the average over the seven items (after reverse coding of Item 2). Other predictors we used are country, gender, education, and work status. In psychological research, the relationship between attitudes and behaviour is often of interest. Therefore, our primary focus is on the relationship between environmental concern and efficacy and the four behavioural items, controlling for the other variables.
Model S Predictors deviance npar AIC BIC
1 1 All 169060.06 39 169138.06 169438.71
2 2 All 164644.90 59 164762.90 165217.74
3 3 All 162513.15 77 162667.15 163260.74
4 3 All - C 173682.04 41 173764.04 174080.11
5 3 All - G 162897.44 74 163045.44 163615.91
6 3 All - E 162644.64 74 162792.64 163363.11
7 3 All - W 162548.75 71 162690.75 163238.09
8 3 All - A 163119.68 74 163267.68 163838.15
9 3 All - EC 163164.41 74 163312.41 163882.87
10 3 All - EE 163487.87 74 163635.87 164206.34
11 3 All - W - C 173786.77 35 173856.77 174126.59
12 3 All - W - G 162938.51 68 163074.51 163598.72
13 3 All - W - E 162678.07 68 162814.07 163338.28
14 3 All - W - A 163378.23 68 163514.23 164038.45
15 3 All - W - EC 163201.58 68 163337.58 163861.79
16 3 All - W - EE 163487.87 74 163635.87 164206.34
Table 1: Model selection for ISSP data from cumulative logistic reduced rank regression. Abbreviations: C: Country, G: Gender, E: Education, W: Work, A: Age, EC: Environmental Concern, EE: Environmental Efficacy.
We start with a predictor matrix $\bm{X}$ of dimension 16696 by 19. First, we need to find an optimal dimensionality (rank). The model with $S=3$ has the lowest AIC and BIC, see Table 1. In the next step, we delete one by one the predictor variables from the model and see whether the AIC and BIC improve. Only for the variable work status the BIC decreases, but the AIC increases. After leaving out working status, we checked whether any other predictor can be left out, which is not the case. Model 7, with 71 parameters is the preferred model.
The estimated reduced rank coefficients (i.e., $\bm{BV}^{\prime}$) from Model 7 are displayed in Table 2. Note that Thailand (TH) serves as a reference country, so all regression coefficients for the countries should be compared to Thailand. For example, the cumulative log odds for Austria (AT) is 2.64 higher for leisure activities outside in nature (OUT), 0.69 lower for eating meet (MEAT), 1.93 higher for special effort to sort glass or tins or plastic or newspapers (RECYCLE), and 0.66 higher for avoiding to buy products for environmental reasons (AVOID). Checking our variables of interest, we see that when environmental concern (EC) and environmental efficacy (EE) increase the cumulative log odds for pro environmental behavior goes up, most strongly for environmental efficacy. For example, with every standard deviation increase in environmental efficacy the cumulative log odds for special effort to sort glass/tins/plastic/newspapers (RECYCLE) goes up by 0.46.
OUT MEAT RECYCLE AVOID
1|2 -1.28 -3.84 -2.50 -1.53
2|3 0.53 -3.26 -0.74 0.95
3|4 2.03 -2.48 0.83 3.38
4|5 3.91 -1.79
5|6 -0.96
6|7 -0.09
7|8 1.22
AT 2.64 -0.69 1.93 0.66
TW 0.45 1.19 1.51 1.02
FI 2.46 -1.02 1.60 0.44
DE 2.75 -0.38 2.18 0.82
HU 1.14 1.66 0.66 0.34
IS 2.28 -1.51 1.48 0.36
JP -0.44 -0.82 1.38 0.95
NZ 2.20 -1.65 1.74 0.54
PH -0.36 0.60 0.52 0.48
RU 0.83 -0.18 -1.24 -1.03
SL 3.19 -0.27 1.73 0.44
CH 2.66 -0.81 2.42 0.97
Female 0.03 0.50 0.22 0.18
Eduyrs 0.03 -0.02 0.19 0.12
age 0.10 0.14 0.40 0.26
EC 0.02 0.10 0.36 0.25
EE 0.19 0.20 0.46 0.28
Table 2: Estimated regression coefficients $\bm{BV}^{\prime}$ for the four pro-environmental behaviours.
4.2 Cumulative Logistic Restricted Multidimensional Unfolding
In this analysis, we like to see whether there are differences between countries concerning environmental efficacy. The seven ordinal items are treated as the responses ($R=7$) and the dummy variables for countries are used as predictors ($P=13-1=12$).
@@ 4.2 Cumulative Logistic Restricted Multidimensional Unfolding
deviance npar AIC BIC
1 316958.70 47.00 317052.70 317415.03
2 315116.83 65.00 315246.83 315747.92
3 314289.43 82.00 314453.43 315085.57
showing that the three-dimensional model fits best.
For illustrative purposes, the two-dimensional biplot is shown in Figure 3. We can discern three groups of countries: on the upper right there are the Asian countries (Taiwan, Japan, Philippines, and Thailand), on the bottom there are the East European countries Russia and Hungary, while on the left hand side there are the more Western countries (Austria, Finland, Germany, Iceland, New Zealand, Slovenia, and Switzerland).
@@
Considering the items, we see that Items 1, 3, 5, and 6 are quite close together, all showing hesitation of personal options (i.e., It is just too difficult for someone like me to do much about the environment; There are more important things to do in life than protect the environment; Many of the claims about environmental threats are exaggerated; I find it hard to know whether the way I live is helpful or harmful to the environment)
Items 4 and 7 are close together but a bit apart from the first cluster (There is no point in doing what I can for the environment unless others do the same; Environmental problems have a direct effect on my everyday life). Item 2 (I do what is right for the environment, even when it costs more money or takes more time) is a bit isolated on the right hand side.
To better visualize the differences between average opinions of people from the different countries, in Figure 4 we show the decision regions for two items (2 and 4). For item 2 (left hand side), the estimated radii corresponding to the thresholds 3|4 and 4|5 are negative, so these circles are not shown. The interpretation is that in none of the countries the average opinion is to agree or strongly agree with this assertion. Therefore, the 13 countries are partitioned into two groups, those who neither disagree nor agree (Finland, Thailand, and the Philippines) and those who disagree (Taiwan, Japan, Russia, Hungary, Austria, Germany, Iceland, New Zealand, Slovenia, and Switzerland). The people from these countries, however, disagree for different reasons as the positions for the countries are on different sides of the item. For Item 4, the estimated radius corresponding to the threshold 4|5 is negative, so this circle is not shown. It can be verified that on average the Taiwanese, Japanese, and Philippine people disagree with this statement, while people from Thailand neither agree nor disagree, and all others countries generally agree with the statement.
@@
Figure 4: Estimated configuration for Environmental Efficacy data with decision regions for Item 2 (a; left) and Item 4 (b; right).
@@ 5 Discussion
In this manuscript, we proposed a novel framework for multidimensionional analysis of ordinal data, where we distinguished between proximity and dominance items. For proximity items, distance representations are used, whereas for dominance items inner product representations are used. A distance representation is obtained using multidimensional unfolding, whereas an inner product representation is obtained by using principal components. When predictor variables are available for the participants, restricted versions of the inner product and distance model are obtained, leading to reduced rank regression and restricted multidimensional unfolding. An Expectation Majorization Minimization (EMM) algorithm is proposed for parameter estimation of the structural part. In the E-step the conditional expectation of the complete data negative log likelihood is taken which is majorized by a least squares function. This least squares function is minimized using well known updating steps. For the estimation of the threshold, standard maximum likelihood methods can be employed, using the structural part as an offset. The algorithm monotonically decreases the negative log likelihood function.
@@
Biplot representations were discussed in detail. For the cumulative logistic PCA and RRR such biplots have also been proposed by Vicente-Villardón and Sánchez, (2014). Our CLPCA biplots are largely the same as those, only the procedure for adding variable markers on the response variable axes differ. Where Vicente-Villardón and Sánchez, (2014) use the posterior probabilities to define the markers, we use the underlying latent variable. The advantage of our idea is that all subsequent markers are represented, whereas the posterior probability of some categories of an outcome variable might never be higher than the probabilities of the rest. Furthermore, our markers directly follow from the model estimates, whereas the posterior probabilities need an extra computational step. The CLMDU biplots are, as far as we know, new. In the CLMDU biplots, the distances between points for the participants and points for the response variables determine the probabilities. The threshold parameters can be included in the biplot as circles. For a response variable with $C_{r}$ categories, $C_{r}-1$ circles are drawn partitioning the space in $C_{r}$ regions.
We applied our models to a large data set from the International Social Survey Programme in Section 4, illustrating the use of cumulative logistic reduced rank regression and cumulative logistic restricted multidimensional unfolding.
In our principal component and reduced rank models a proportional odds assumption is made. In the current context, it is not easy to test the validity of this assumption. For the reduced rank regression parameterisation, one possible solution is to fit separate proportional odds regression models for each response variable and checking the assumption. If validated for each response variable, we can safely analyze the variables together in our reduced rank model.
We focused on predictor variables describing the participants. In some data sets, external information about the items is available. In our analysis, we could linearly constraint the matrices $\bm{V}$ to include such information. For the type of response variables considered in our applications, however, typically no further information is available.
@@
Logistic multidimensional models are closely related to Item Response Models. In fact, the model formulation of cumulative logistic PCA in 1 dimension is equivalent to Samejima’s Graded Response Model (GRM; Samejima,, 1969). In the GRM, however, the principal scores are usually assumed to be a random effects from a normal distribution. In PCA, principal scores are fixed effects. Also explanatory versions of the GRM have been proposed (Tuerlinckx and Wang,, 2004) to include predictor variables. Within the context or item response theory (IRT) also single peaked items can be modelled. The best known model for ordinal data is the generalized graded unfolding model (GGUM; Roberts et al.,, 2000). The model definition is quite involved, defined by two GRMs, one from below and one from above. The GGUM model is a unidimensional model, no multidimensional generalizations have been proposed so far, although recently a R-package for estimation of such multidimensional models has been proposed (Tu et al.,, 2021). Recently some explanatory versions, including predictors for the observations have been proposed by Usami, (2011) and Joo et al., (2022). Maximum likelihood estimation of logistic models with normally distributed random effects is generally problematic due to an intractable integral (Tuerlinckx et al.,, 2006). In our logistic multidimensional data analysis techniques we do not have those random effects, so maximum likelihood estimation is simpler. The goals of the two approaches are different. Item response models usually are targeted towards optimal latent trait estimation. Often, a priori knowledge is available on the traits under investigation. External information (i.e., predictor variables) is used to address subpopulation heterogeneity or increase estimation accuracy. The goals of our analysis framework is more towards dimension reduction to obtain insight into the structure of the response variables or, when predictor variables are available, to develop simultaneous regression models for the response variables in a reduced dimensional space.
@@
MdR initiated and conceptualized this study. MdR developed the algorithm and implemented it in R. FB wrote C-code for multidimensional unfolding and restricted multidimensional unfolding. DW tested the CLMDU and CLRMDU algorithms and applied them to the ISSP data. DW developed the visualization of MDU and RMDU analysis under the supervision of MdR. LB tested the CLPCA and CLRRR algorithms and applied them to the ISSP data. MdR wrote the first draft of the paper. FB, LB, and DW commented on the first draft. All authors approved the final manuscript.

### v2

@@ 4.3 ISSP data: Environmental Efficacy
Participants had to indicate on a five-point Likert scale for each of these statements whether they agreed strongly (coded as 5), agreed (4), are neutral (3), disagreed (2), or disagreed strongly (coded as 1).
As predictor variables, we use the 12 zero-one coded dummy variables for the countries, using Thailand as the control category (i.e., coded with 12 zeros), a dummy variable for gender where female is coded 1 and therefore male serves as a reference, and both the standardised scores of age and number of years of education as numeric predictors. The origin of the Euclidean space, therefore, corresponds to male participants from Thailand having average age and education.
@@
Dimensionality deviance npar aic bic
1 316958.57 50 317058.57 317444.02
2 313685.65 71 313827.65 314374.99
3 312597.80 91 312779.80 313481.32
showing that the three-dimensional model fits best. For illustrative purposes, the two-dimensional biplot is shown in Figure 5. The response variables are located from the bottom left (item 1, too difficult) to top middle (item 2, do right). They lie almost on a curve, where items 5 and 6 (exaggerated and hard to know) lie close together, indicating similar response tendencies for those two items.
@@
Looking at the predictor side we first note that the origin, where the variable axes for Education and Age cross represents male participants from Thailand with average age and education. We can see that Education has a large influence on the positioning of the participants, where participants with a few years of education are in the top of the biplot and those with many years of education in the bottom. The variable axis for Age is much shorter, and therefore Age has a smaller effect on the outcomes. The last control variable is Gender, where we see that the category Female is below the origin. Therefore, comparing female and male participants with the same values for the other predictor variables, the females are located below male participants and therefore closer to the items at the bottom of the configuration. The countries partition in two clusters, with on the right hand side of the biplot the Asian countries Japan, the Philippines, Taiwan and Thailand and on the left hand side a cluster of Hungary, Russia, New Zealand, Island, Slovenia, Switzerland, Austria, Germany, and Finland. Austria and Germany are very close, meaning that responses in these two countries follow a similar pattern. Russia and Hungary are close together and positioned more in the top of the biplot, further away from the response variables more important, exaggerated, hard to know, and too difficult, indicating lower probabilities of agreement with these items compared to participants from other countries.
Now, let us look in more detail at two biplots where we included the circles representing the classification regions for items 4 (“There is no point in doing what I can for the environment unless others do the same”) and 7 (“Environmental problems have a direct effect on my everyday life”). The biplots are shown in Figure 6. For item 4, we see that the cluster of Asian countries on the left disagrees (all participants from Taiwan and most participants from the Philippines) or is neutral (most participants from Thailand and the higher educated and elderly in Japan).
@@
Figure 6: Estimated configuration for Environmental Efficacy data with decision regions for Item 4 (a) and Item 7 (b).
Participants from the other countries mainly agree with this statement, because their position is within the neutral | agree (i.e., 3|4) circle. An exception are the participants with only few years of education from Russia and Hungary, who tend to be neutral.
Inspecting the biplot with classification regions for item 7, we can conclude that only participants from Finland with average age and education tend to agree with this item. Most participants from the Asian countries tend to disagree (except participants from Thailand), while most participants from the other countries tend to be neutral towards this item. It seems that especially Age has an influence on this response variable as elderly people from the Asian countries have larger probabilities to agree, while the younger people from the other cluster have higher probabilities to agree (i.e., they are closer to the position of the item).
5 Simulation studies
We conducted a simulation study to verify whether the algorithms work properly. Therefore, we use the estimated parameters from the second example, described in Section 4.2, specifically the matrices $\bm{B}$ and $\bm{V}$. The precise values are shown in Appendix C. In the data generation process, we start drawing predictor variables from the multivariate normal distribution with mean zero and covariance matrix equal to the correlation matrix of the predictors in Section 4.2 (see Appendix C for values). The predictor variables and the population matrix $\bm{B}$ defines $\bm{U}$. With these coordinates and the matrix $\bm{V}$, values on the latent variables (i.e., the $\theta_{ir}$’s) can be computed. For the dominance models we use the inner product ($\theta_{ir}=\langle\bm{B}^{\prime}\bm{x}_{i},\bm{v}_{r}\rangle$), while for the proximity models we use the distances ($\theta_{ir}=-d(\bm{B}^{\prime}\bm{x}_{i},\bm{v}_{r})$). With these values and the threshold parameters, probabilities of the response categories for each of the response variables can be obtained. We draw observed outcome variables for each response variable independently from the multinomial distribution.
We vary sample size with values 250, 500, and 1000. We also vary whether the response variables have three or five categories and we manipulate the number of response variables, to be equal to 4 or 8. Note that in the empirical example we had 4 response variables. We created 4 others by simply rotating $\bm{V}$ by $45^{o}$ and adding the obtained coordinates to the matrix (see Appendix C for population values). We used a full factorial $3\times 2\times 2$ design, with 200 replications per condition.
The simulation is done separately for the cumulative logistic reduced rank model and the cumulative logistic restricted multidimensional unfolding.
As outcome variable we take the following measure of recovery:
$\delta=\sqrt{\frac{\sum_{i}\sum_{r}(\theta_{ir}-\hat{\theta}_{ir})^{2}}{\sum_{i}\sum_{r}\hat{\theta}^{2}_{ir}}},$
where $\theta_{ir}=\langle\bm{B}^{\prime}\bm{x}_{i},\bm{v}_{r}\rangle$ for the dominance model and $\theta_{ir}=-d(\bm{B}^{\prime}\bm{x}_{i},\bm{v}_{r})$ for the proximity model. This measure resembles the Stress-1 value that is often used in multidimensional scaling and unfolding. The benefit of this measure over, for example, looking at the recovery of $\bm{B}$ or $\bm{V}$ is that we do not need to take into account any reflection, rotation, or other transformations to overcome indeterminacies. We like to point out that the measure $\delta$ is not necessarily comparable between the dominance and proximity perspective as for the proximity models the values of $\theta_{ir}$ are all negative, while for the dominance models they can be both positive and negative.
Figure 7: Simulation results for cumulative logistic reduced rank regression. $R$ denotes the number of response variables, $C$ the number of response categories per response variable. On the horizontal axis we show the different sample sizes, while on the vertical axes, the value of recovery is found where lower values represent better recovery.
The results of our simulation studies are presented in Figures 7 for the reduced rank model and Figure 8 for the restricted multidimensional unfolding. In both cases, we see that recovery is good and that it improves with increasing sample size, number of response variables ($R$), and number of response categories ($C$). The changes for the number of response variables and the number of response categories are small, but the values of $\delta$ become smaller and less variable. Sample size has a larger influence as is clear from the boxplots.
Figure 8: Simulation results for cumulative logistic restricted multidimensional unfolding. $R$ denotes the number of response variables, $C$ the number of response categories per response variable. On the horizontal axis we show the different sample sizes, while on the vertical axes, the value of recovery is found where lower values represent better recovery.
@@ 6 Discussion
In this manuscript, we proposed a novel framework for multidimensional analysis of ordinal data in a cumulative logistic framework. Logistic regression models are the usual models of choice when the response variable are categorical. For an ordinal response variable, the cumulative logistic regression model, also known as the proportional odds regression model, is a typical analysis model often used in practice. We developed cumulative logistic models for multiple ordinal response variables that we like to analyze together.
We distinguished between models for proximity and dominance items. For proximity items, distance representations are used, whereas for dominance items inner product representations are used. A distance representation is obtained using multidimensional unfolding, whereas an inner product representation is obtained using principal components. When predictor variables are available for the participants, restricted versions of both types of models are obtained, leading to reduced rank regression and restricted multidimensional unfolding.
For maximum likelihood estimation, an Expectation Majorization Minimization (EMM) algorithm was developed and tested. In the E-step the conditional expectation of the complete data negative log likelihood is taken which is majorized by a least squares function. This least squares function is minimized using well known updating steps. For the estimation of the threshold, standard maximum likelihood methods can be employed, using the structural part as an offset. The algorithm monotonically decreases the negative log likelihood function. In a simulation study, we tested the performance of the algorithm and we showed that overall the recovery is good. In more detail, the recovery increases with sample size, more response variables, and increasing number of response categories.
@@
Biplot representations were discussed in detail. For cumulative logistic PCA similar biplots have been proposed before by Vicente-Villardón and Sánchez, (2014). Our CLPCA biplots differ in the way the markers for the response variable axes are defined. We use the underlying latent variable, whereas Vicente-Villardón and Sánchez, (2014) use posterior probabilities. The advantage of our idea is that all subsequent markers are represented, whereas the posterior probability of some categories of an outcome variable might never be higher than the probabilities of the rest. Furthermore, our markers directly follow from the model estimates, whereas the posterior probabilities need an extra computational step. The CLMDU biplots are, as far as we know, new. In the CLMDU biplots, the distances between points for the participants and points for the response variables determine the probabilities. The threshold parameters can be included in the biplot as circles. For a response variable with $C_{r}$ categories, $C_{r}-1$ circles are drawn partitioning the space in $C_{r}$ regions.
We showed applications of the analysis framework for three empirical data sets. The response variables in the first data set are ten exam items about the application of statistics. These items are typical cognitive response variable and therefore are of the dominance type. Besides the responses to ten items also characteristics of the students are available. We used these characteristics as predictor variables. For the analysis of these data, we employed the cumulative logistic reduced rank model and show the results both in terms of a table with regression coefficients (Table A2) and a biplot (Figure 3). Each of the columns of the coefficient matrix can be interpreted in a similar way as a standard proportional odds model. Although this allows for a detailed interpretation of the effect of each predictor on each response, it is difficult to obtain a holistic interpretation of the model or a view about the dependencies among the response variables from this table. The biplot enhances these two latter aspects. We saw that a group of eight response variables are in a similar way related to the predictor variables, the variable axes for these response variables all run from left to right. Two response variables behave differently, running from the bottom to the top (A2g and A1d). We can infer that there are strong dependencies among the group of eight and the group of 2. Furthermore, we can obtain a holistic view on the model by looking at the angles of the predictor and response variable axes. As there is a group of predictor variables with axes running from top left to bottom right (APS, SATSv, SATSc, ENGc) and another from left to right (SATSa, PMP, ENGb) we can easily remember that those two groups act differently on the response variables. The variables within either of the two groups act similarly on the response variables. Specific effects can be obtained from the table again. We see that the table and biplot reinforce each other in interpretation.
In the second data set there are four behavioural response items. Predictor variables include gender, age, education, and environmental efficacy and concern. We used this data set to contrast the distance and inner product approach. The data were analysed by both approaches and it was shown that the distance approach fits the data better, mainly because of one of the response variables. In the biplot visualization, the other three response variables are represented by a point in the periphery of the plot indicating that a dominance perspective suits these three items well. The response variable OUT, however, acts like a proximity item: participants answering that they never engage in outdoor activity have different characteristics in terms of age, number of years of education, environmental concern and environmental efficacy.
Using the third data set, we applied the distance approach. The response variables measure attitudes which usually are considered proximity variables. Given age, gender, and education, the goal was to see whether there are differences between countries. We showed the resulting biplot visualization and gave a detailed interpretation indicating the differences between countries and the influence of education, age and gender. The application of our framework to these data sets shows the versatility of our approach.
The framework we proposed gives the opportunity to analyse complex data sets having ordinal response variables with dimension reduction techniques. The response variables do not need to be indicators of underlying constructs as in psychological scales but simply may be a number of related questions on a given topic (like pro environmental behaviour). The framework is applicable even in the case of a limited number of response variables, in contrast to item response models that usually require many items (see below for a further comparison). The models for proximity items and the corresponding visualizations with biplots using circles as classification regions are new and are a valuable tool for data analysis. Biplots for ordinal response variables have been proposed by (Vicente-Villardón and Sánchez,, 2014). They, however, only show methods for dominance variables without predictor variables. We extended their method by constraining the principal scores using predictor variables (i.e., the reduced rank regression approach). We also gave an alternative way for adding markers to the variable axes, one that is in line with the underlying variable idea.
In our multidimensional models a proportional odds assumption is made, as discussed in Section 2.1. It is not easy to test the validity of this assumption within our modelling approach. One approach would be to fit different models that do not make this assumption and compare the fit of the two models. For the CLPCA model, we could compare against multinomial multiple correspondence analysis (MMCA; Groenen and Josse,, 2016). For CLRRR we could compare to a constraint version of MMCA (which first needs to be developed). Such comparisons are related to the score test for the proportional odds assumption (Agresti,, 2002, Section 7.2). This score test tests whether the effects are the same for each cumulative logit against the alternative of separate effects. If the fit of the MMCA models is much better this is an indication of a violation of the proportional odds assumption for one or more response variables. For which response variable the assumption is violated then needs further investigation. An alternative applicable for CLRRR, is to apply the score test for each response variable separately when fitting $R$ different models. When none of these tests suggest a rejection of the null hypothesis, we can safely conclude that the assumption is also valid for the multivariate model. When one of the tests rejects the null hypothesis, we directly know for which response variable the proportional odds assumption is false. For the distance models (CLMDU and CLRMDU) validating the proportional odds assumption is more difficult. We could, for instance, develop new models that combine the cumulative logit models developed here with multinomial restricted unfolding (De Rooij and Busing,, 2024), a distance model for nominal variables, and investigate whether such a new model signals violations of the assumption. Further work is needed to find ways to verify this assumption within the framework presented.
On the other hand, Harrell, (2001) points out that we should not worry too much about the proportional odds assumption. Rank order based statistics such as the Wilcoxon test and the Kruskal-Wallis tests are special cases of the proportional odds regression model. Furthermore, rank order correlations are closely related to the proportional odds model. It seems best, to place our models in a bias-variance trade-off perspective. The proportional odds assumption might not be valid for some response variables leading to biased results. However, adding parameters to avoid the assumption might lead to more variance and, when the extra variance exceeds the bias, lead to worse model performance.
Sometimes, a priori information is available about which response variables group together on specific dimensions. This would entail that elements of the matrix $\bm{V}$ are set to zero. Similar constraints could, in theory, be imposed on the matrix $\bm{B}$, specifying that some predictor variables are connected to specific dimensions. At the moment it is not possible to use such information in the analysis. Further research is needed to incorporate such knowledge.
@@
Logistic multidimensional models are closely related to Item Response Models. In fact, the model formulation of cumulative logistic PCA in one dimension is equivalent to Samejima’s Graded Response Model (GRM; Samejima,, 1969). In the GRM, however, the principal scores are usually assumed to be a random effects from a normal distribution. In PCA, principal scores are fixed effects. Also explanatory versions of the GRM have been proposed (Tuerlinckx and Wang,, 2004) to include predictor variables. Item response models are usually targeted towards test construction. Our models focus more on dimension reduction and prediction of correlated response variables. We like to note that these graded response models also make the proportional odds assumption, similar to our approaches, although not much emphasis is placed on this assumption in the literature. Within the context or item response theory, also single peaked items can be modelled. The best known model for ordinal data is the generalized graded unfolding model (GGUM; Roberts et al.,, 2000). The model definition is quite involved, defined by two GRMs, one from below and one from above. The GGUM model is a unidimensional model, no multidimensional generalizations have been proposed so far, although recently a R-package for estimation of such multidimensional models has been proposed (Tu et al.,, 2021). Recently some explanatory versions, including predictors for the observations have been proposed by Usami, (2011) and Joo et al., (2022). Again, those item response models are targeted towards test construction, whereas our approaches are targeted towards dimension reduction and prediction of correlated response variables. Maximum likelihood estimation of logistic models with normally distributed random effects is generally problematic due to an intractable integral (Tuerlinckx et al.,, 2006). In our logistic multidimensional data analysis techniques, we do not have those random effects, so maximum likelihood estimation is simpler. The goals of the two approaches are different. Item response models usually are targeted towards optimal latent trait estimation. Often, a priori knowledge is available on the traits under investigation. External information (i.e., predictor variables) is used to address sub population heterogeneity or to increase estimation accuracy. The goals of our analysis framework is more towards dimension reduction to obtain insight into the structure of the response variables or, when predictor variables are available, to develop simultaneous regression models for the response variables in a reduced dimensional space.
In our framework, we focused on predictor variables describing the participants. In some situations, external information about the items might be available. In our analysis, we could linearly constrain the matrices $\bm{V}$ to include such information. For the type of response variables considered in our applications, however, typically no further information is available. Further research and programming is needed to incorporate such constraints.
In conclusion, we proposed a family of models for multidimensional analysis of multiple ordinal response variables. The framework contains four different models. We distinguished between models for dominance variables and proximity variables. Within each we distinguished models with and without predictor variables. Algorithms for all methods proposed in this paper are implemented in the R software. The logistic mapping package (De Rooij et al.,, 2024) contains the functions clpca and clmdu and corresponding plotting function that can be used for the analysis described in this paper.
@@
Thanks:
The authors would like to thank Dr. Rosa Fabbricatore for making the data available to us. We would also like to thank the three anonymous reviewers for their helpful input.
The first author revised the manuscript while he was a fellow at the Netherlands Institute of Advanced Studies in Amsterdam.
@@
The data set about Students performance on the statistics exam can be requested from Dr. Rosa Fabbricatore
@@
MdR initiated and conceptualized this study. MdR developed the algorithm and implemented it in R. FB wrote C-code for multidimensional unfolding and restricted multidimensional unfolding. DW tested the CLMDU and CLRMDU algorithms and applied them to the ISSP data. DW developed the visualization of MDU and RMDU analysis under the supervision of MdR. LB tested the CLPCA and CLRRR algorithms and applied them to the ISSP data. MdR wrote the first draft of the paper. FB, LB, and DW commented on the first draft. MdR revised the manuscript on the basis of th reviewers comments. All authors approved the final manuscript.
Appendix A: Statistics for Student data
We use different labels for the response variables than the authors in Fabbricatore et al., (2024) use. Here is a table relating our labels (left column) to the names shown in their paper (center column) and the actual question for the students:
A1a T1_ClassVar_A Select the type of variable that best describes the gross annual income
A1b T1_GraphForQuant_A Find the error in a graph representing the distribution of a continuous variable (i.e. waiting time in minutes)
A1c T1_Median_A Calculate the median of individual series data
A1d T1_ArithmeticMean_A Calculate the overall weighted arithmetic mean given the mean of three groups of individuals with different class sizes
A1g T1_SkewnessNKurtosis_A Calculate the Hotelling-Solomon skewness coefficient
A2a T2_SamplingSpace_A Calculate the size of the sample space if 1 coin and 2 six-sided dice are tossed together at the same time
A2c T2_ConditionalProb_A Calculate the conditional probability, given the marginal probabilities and the joint probability
A2d T2_Bernoulli_A Calculate the variance of a Bernoulli trial given the probability of success
A2g T2_Gaussian_A Calculate probability for normally distributed data
A2i T2_SamplingVariance_A Determine the sample size using the formula derived from the corrected sample variance
The model selection procedure is summarized using Table A1.
npar deviance aic bic
1 48.00 2310.47 2406.47 2546.98
2 74.00 2250.50 2398.50 2615.12
3 98.00 2206.37 2402.37 2689.24
Age 72.00 2251.38 2395.38 2606.14
Gender 72.00 2255.71 2399.71 2610.48
Math 72.00 2277.05 2421.05 2631.82
Statistical.Anxiety 68.00 2253.18 2389.18 2588.23
SATS 66.00 2271.74 2403.74 2596.94
MSLQ 66.00 2257.26 2389.26 2582.46
APS 72.00 2258.28 2402.28 2613.04
AMS 72.00 2250.55 2394.55 2605.31
ENG 68.00 2265.35 2401.35 2600.40
Table A1: Fit statistics for the student data. First three rows show fit statistics of models in 1 to 3 dimensions including all predictors. The remaining rows show fit statistics leaving out one variable or (set of) scales. npar denotes the number of parameters.
The implied coefficients ($\bm{BV}^{\prime}$) are shown in Table A2, and can simply be understood similarly as coefficients in a cumulative logistic model (proportional odds model).
A1a A1b A1c A1d A1g A2a A2c A2d A2g A2i
0||||1 -0.85 -2.98 -4.03 -1.38 -2.53 -2.57 -0.88 -2.03 -1.26 -0.92
1||||2 1.22 0.30 -0.45 -0.34 -1.78 -2.07 0.42 -1.79 0.06 1.18
G -0.03 -0.61 -0.19 -0.00 -0.17 -0.68 -0.30 -0.61 0.27 -0.96
PMP 0.05 0.37 0.29 0.38 0.23 0.55 0.27 0.54 0.45 0.47
SATSa 0.02 0.11 0.14 0.23 0.10 0.20 0.11 0.21 0.31 0.11
SATSc 0.01 -0.26 0.05 0.28 0.02 -0.20 -0.06 -0.14 0.56 -0.49
SATSv 0.01 -0.20 0.02 0.17 -0.00 -0.17 -0.06 -0.13 0.36 -0.37
SATSd -0.02 -0.06 -0.13 -0.23 -0.09 -0.14 -0.08 -0.16 -0.34 -0.02
APS 0.00 -0.25 0.00 0.17 -0.02 -0.23 -0.09 -0.18 0.38 -0.45
ENGa 0.01 0.02 0.07 0.13 0.05 0.07 0.04 0.08 0.20 -0.00
ENGb 0.00 0.14 0.03 -0.02 0.03 0.15 0.06 0.13 -0.09 0.23
ENGc -0.01 0.21 -0.06 -0.26 -0.02 0.15 0.05 0.10 -0.51 0.41
Table A2: Implied coefficient for the student data
Appendix B: Pro-environment behaviour
The biplots for the cumulative logistic restricted multidimensional unfolding analysis including the circles for each of the response variables is shown in Figure B1.
Figure B1: Biplot for the cumulative logistic restricted multidimensinal unfolding solution relating environmental attitudes with pro-environmental behaviour.
The biplot for the cumulative logistic reduced rank model is shown in Figure B2.
Figure B2: Biplot for the cumulative logistic reduced rank model relating environmental attitudes with pro-environmental behaviour.
Appendix C: Population parameters for Simulation studies
The predictor variables are sampled from a multivariate normal distribution with zero means and covariance matrix
X1 X2 X3 X4 X5
X1 1.00 0.01 -0.02 0.01 0.04
X2 0.01 1.00 -0.59 0.19 0.16
X3 -0.02 -0.59 1.00 -0.00 -0.00
X4 0.01 0.19 -0.00 1.00 0.25
X5 0.04 0.16 -0.00 0.25 1.00
The matrix with population coefficients $\bm{B}$ equals the estimated coefficients from the example in Section 4.2, that is
1 2
X1 -0.16 0.19
X2 -0.37 0.04
X3 -0.17 0.19
X4 -0.40 -0.17
X5 -0.28 0.12
The population matrix $\bm{V}$ equals
1 2
Y1 0.44 -0.45
Y2 0.34 2.35
Y3 -1.68 0.05
Y4 -1.55 0.08
Y5 -0.16 -0.61
Y6 2.18 0.94
Y7 -0.84 1.45
Y8 -0.74 1.36
where we use only the first four rows when $R=4$ and the complete matrix when $R=8$
The parameters $\bm{m}$ are equal for the different response variables. The threshold values in case of three categories are -1.0 and -0.5, while the values with five response categories are -2.0, -1.5, -1.0, and -0.5.